# Stochastic Optimal Control of an Industrial Power-to-Heat System with High-Temperature Heat Pump and Thermal Energy Storage

### v2

@@ Appendix C Model Parameters
$c_{\text{S}}$ Storage specific heat capacity 1.025 [kJ/kgK]
$c_{\text{F}}$ Thermal oil specific heat capacity 2.314 [kJ/kgK]
@@
${\lambda}_{W}$ Estimate for mean-reversion speed of $Y^{W}$ 0.1702 [1/$s$]
${\sigma}_{W}$ Estimate for volatility of $Y^{W}$ 0.2486 [m/$\sqrt{\text{s}^{3}}$]
@@
${\lambda}_{S}$ Estimate for mean-reversion speed of $Y^{S}$ 0.2534 [1/$s$]
@@
(b) Parameters for the numerical solution of the MDP with backward dynamic programming and Q-learning.
@@
$n_{A}$ Number of grid-points for control $A$ $31$ [-]
@@
$k^{Ref}$ $\sigma$-rule parameter for $\pazocal{X}^{Ref}_{n}$ $3$ [-]
$k^{Ext}$ $\sigma$-rule parameter for $\pazocal{X}_{n}$ $4$ [-]
@@
$\varepsilon_{0}$ Exploration rate linearly decaying to zero $1$ [-]
@@
Parameters for the Terminal Condition Value Unit
@@
$s_{\text{Liq}}$ Liquidation price 00 [€/Mwh]
@@ Appendix D Details on Thermal Energy Storage Operational Modes
Due to $T^{\text{HT,out}}_{n}>T^{\text{SG,in}}$ and $l^{\text{C}}\in(0,1]$ it follows directly $T^{\text{C}}(t)<T^{\text{SG,in}}$, so that the charging outlet temperature never exceeds the SG inlet temperature. By charging the TES, its temperature $R$ increases linearly over time (3) and so, based on (84), $T^{\text{C}}$ also changes with time. For this reason, the charging factor must be adjusted over time to ensure a constant SG inlet temperature. Combining (8), (84) and (85) leads to the following relation for the charging factor
@@
which must hold in each period $[t_{n},t_{n+1})$ and describes the dependence of $l^{\text{C}}$ on the chosen control $A_{n}$ and the TES temperature $R$.
@@
Analogous to (85), the discharging outlet temperature $T^{\text{D}}$ is given by
@@
As above, the discharging factor must also be continuously adjusted in each period to ensure the constant HTHX inlet temperature $T^{\text{HT,in}}_{n}$. Substituting (87) into (88) and using (8) yields for $t\in[t_{n},t_{n+1})$
@@
Note that the time-varying charging and discharging factors (86) and (89) depending on $R$ and chosen $A_{n}$ are used to derive the control conditions in Appendix G.
Appendix E Proofs
E.1 Proof of Lemma 3.1
@@ Appendix E Proofs
First let’s consider the Ornstein-Uhlenbeck process $Y^{W}(t)$ and find a solution on $[t_{a},t_{b}]$ with initial condition $Y^{W}(t_{a})=y_{W}\in\mathbb{R}$. Applying Itô’s formula with $F(t,Y^{W}(t))=e^{\lambda_{W}t}Y^{W}(t)$ yields
@@
with $I_{W}(t)=\int_{t_{a}}^{t}\sigma_{W}e^{-\lambda_{W}(t-r)}\mathrm{d}B^{W}(r)$. The stochastic integral appearing in (91) is a mean-zero martingale. Using this fact the conditional expected value of $Y^{W}(t)$ given $Y^{W}(t_{a})=y_{W}$ reads
@@
with $\tau=t-t_{a}$. The variance follows from the Itô isometry leading to
@@
In order to obtain a solution for $Y^{S}(t)$ Itô’s formula with $F(t,Y^{S}(t))=e^{\lambda_{S}t}Y^{S}(t)$ is applied leading to
@@
Again using the integral form with initial condition $Y^{S}(t_{a})=y_{S}$ on $[t_{a},t_{b}]$ and multiplying by $e^{-\lambda_{S}t}$ yields
@@
with $I_{S,1}(t)=\int_{t_{a}}^{t}\sigma_{S}e^{-\lambda_{S}(t-r)}\mathrm{d}B^{S}(r)$ and $I_{S,2}(t)=-\lambda_{S}c_{W}\int_{t_{a}}^{t}e^{-\lambda_{S}(t-r)}Y^{W}(r)\mathrm{d}r$.
The integral expression $I_{S,2}(t)$ can further be simplified by using the solution for $Y^{W}(t)$, see (91),
$I_{S,2}(t)$ $=-\lambda_{S}c_{W}\int_{t_{a}}^{t}e^{-\lambda_{S}(t-r)}\bigg{(}e^{-\lambda_{W}(r-t_{a})}y_{W}+\int_{t_{a}}^{r}\sigma_{W}e^{-\lambda_{W}(r-u)}\mathrm{d}B^{W}(u)\bigg{)}\mathrm{d}r$ (96)
@@
with $I_{S,2.1}(t)=\int_{t_{a}}^{t}e^{-\lambda_{S}(t-r)}e^{-\lambda_{W}(r-t_{a})}y_{W}\mathrm{d}r$ and $I_{S,2.2}(t)=\int_{t_{n}}^{t}\int_{t_{a}}^{r}\sigma_{W}e^{-\lambda_{S}(t-r)}e^{-\lambda_{W}(r-u)}\mathrm{d}B^{W}(u)\,\mathrm{d}r$.
Evaluation each integral individually gives
$I_{S,2.1}(t)$ $=e^{-\lambda_{S}t}e^{-\lambda_{W}t_{a}}y_{W}\int_{t_{a}}^{t}e^{(\lambda_{S}-\lambda_{W})r}\mathrm{d}r=\frac{e^{-\lambda_{S}t}e^{-\lambda_{W}t_{a}}y_{W}}{\lambda_{S}-\lambda_{W}}(e^{(\lambda_{S}-\lambda_{W})t}-e^{(\lambda_{S}-\lambda_{W})t_{a}})$ (98)
$=\frac{y_{W}}{\lambda_{S}-\lambda_{W}}(e^{-\lambda_{W}(t-t_{a})}-e^{-\lambda_{S}(t-t_{a})})$ (99)
and for $I_{S,2.2}(t)$ by interchanging the order of integration
$I_{S,2.2}(t)$ $=\int_{t_{a}}^{t}\int_{u}^{t}\sigma_{W}e^{-\lambda_{S}(t-r)}e^{-\lambda_{W}(r-u)}\mathrm{d}r\,\mathrm{d}B^{W}(u)=\sigma_{W}e^{-\lambda_{S}t}\int_{t_{a}}^{t}e^{\lambda_{W}u}\int_{u}^{t}e^{(\lambda_{S}-\lambda_{W})r}\mathrm{d}r\,\mathrm{d}B^{W}(u)$ (100)
$=\frac{\sigma_{W}e^{-\lambda_{S}t}}{\lambda_{S}-\lambda_{W}}\int_{t_{a}}^{t}e^{\lambda_{W}u}(e^{(\lambda_{S}-\lambda_{W})t}-e^{(\lambda_{S}-\lambda_{W})u})\mathrm{d}B^{W}(u)$ (101)
@@
$\begin{split}Y^{S}(t)=&y_{S}e^{-\lambda_{S}(t-t_{a})}+\int_{t_{a}}^{t}\sigma_{S}e^{-\lambda_{S}(t-r)}\mathrm{d}B^{S}(r)\\ &-\frac{\lambda_{S}c_{W}}{\lambda_{S}-\lambda_{W}}\bigg{(}y_{W}(e^{-\lambda_{W}(t-t_{a})}-e^{-\lambda_{S}(t-t_{a})})+\sigma_{W}\int_{t_{a}}^{t}(e^{-\lambda_{W}(t-r)}-e^{-\lambda_{S}(t-r)})\mathrm{d}B^{W}(r)\bigg{)}.\end{split}$ (103)
Analogous to the calculation for $Y^{W}(t)$ it is pointed out that both stochastic integrals are martingales with mean zero Gaussian processes. Furthermore, for the expected value of $Y^{S}(t)$ we get
@@
with $\tau=t-t_{a}$. For the variance one only has to consider the variances of the stochastic integrals, due to the independence of $B^{W}$ and $B^{S}$, resulting in
$\textrm{Var}(Y^{S}(t))$ $=\mathbb{E}\bigg{[}\bigg{(}\int_{t_{a}}^{t}\sigma_{S}e^{-\lambda_{S}(t-r)}\mathrm{d}B^{S}(r)\bigg{)}^{2}\bigg{]}+\mathbb{E}\bigg{[}\bigg{(}\int_{t_{a}}^{t}\tfrac{\lambda_{S}c_{W}\sigma_{W}}{\lambda_{S}-\lambda_{W}}(e^{-\lambda_{W}(t-r)}-e^{-\lambda_{S}(t-r)})\mathrm{d}B^{W}(r)\bigg{)}^{2}\bigg{]}$ (105)
@@
The last equality follows from the Itô isometry. Note that the first expected value is derived the same way as for $Y^{W}(t)$, i.e.
@@
The second expected value is calculated by multiplying $(e^{-\lambda_{W}(t-r)}-e^{-\lambda_{S}(t-r)})^{2}$, resulting in
$\mathbb{E}[I_{S,2}(t)^{2}]$ $=\bigg{(}\tfrac{\lambda_{S}c_{W}\sigma_{W}}{\lambda_{S}-\lambda_{W}}\bigg{)}^{2}\mathbb{E}\bigg{[}\int_{t_{a}}^{t}\big{(}e^{-2\lambda_{W}(t-r)}-2e^{-\lambda_{W}(t-r)}e^{-\lambda_{S}(t-r)}+e^{-2\lambda_{S}(t-r)}\big{)}\mathrm{d}r\bigg{]}$ (108)
$=\bigg{(}\tfrac{\lambda_{S}c_{W}\sigma_{W}}{\lambda_{S}-\lambda_{W}}\bigg{)}^{2}\bigg{(}\int_{t_{a}}^{t}e^{-2\lambda_{W}(t-r)}\mathrm{d}r-2\int_{t_{a}}^{t}e^{-(\lambda_{S}+\lambda_{W})(t-r)}\mathrm{d}r+\int_{t_{a}}^{t}e^{-2\lambda_{S}(t-r)}\mathrm{d}r\bigg{)}$ (109)
$=\bigg{(}\tfrac{\lambda_{S}c_{W}\sigma_{W}}{\lambda_{S}-\lambda_{W}}\bigg{)}^{2}\bigg{(}\frac{\Sigma_{Y^{W}}^{2}(\tau)}{\sigma_{W}^{2}}+\frac{\Sigma_{Y^{S}}^{2}(\tau)}{\sigma_{S}^{2}}-\frac{2}{\lambda_{S}+\lambda_{W}}\big{(}1-e^{-(\lambda_{S}+\lambda_{W})\tau}\big{)}\bigg{)}$ (110)
@@
In the third equality we used $\int_{t_{a}}^{t}e^{-2\lambda_{W}(t-r)}\mathrm{d}r=\nicefrac{{\Sigma_{Y^{W}}^{2}(\tau)}}{{\sigma_{W}^{2}}}$ and $\int_{t_{a}}^{t}e^{-2\lambda_{S}(t-r)}\mathrm{d}r=\nicefrac{{\Sigma_{Y^{S}}^{2}(\tau)}}{{\sigma_{S}^{2}}}$.
Summarizing the variance of $Y^{S}(t)$ is given by
$\Sigma_{S}^{2}(\tau)=\textrm{Var}(Y^{S}(t))=\Sigma_{Y^{S}}^{2}(\tau)+\bigg{(}\tfrac{\lambda_{S}c_{W}\sigma_{W}}{\lambda_{S}-\lambda_{W}}\bigg{)}^{2}\bigg{(}\Sigma_{Y^{W}}^{2}(\tau)+\frac{\sigma_{W}^{2}}{\sigma_{S}^{2}}\Sigma_{Y^{S}}^{2}(\tau)-\frac{2\sigma_{W}^{2}}{\lambda_{S}+\lambda_{W}}(1-e^{-(\lambda_{S}+\lambda_{W})\tau})\bigg{)}.$ (112)
Again since $Y^{S}(t)$ is a sum of normal distributed random variables is holds that
@@
In order to show that $Y^{W}(t)$ and $Y^{S}(t)$ define a bivariate Gaussian random variable, one has to shown that $aY^{W}(t)+bY^{S}(t)$ is normally distributed for all $a,b\in\mathbb{R}$. Let $a,b\in\mathbb{R}$. Let $t\in[t_{a},t_{b}]$ then by the closed-form solutions given $Y^{W}(t_{a})=y_{W}$ and $Y^{S}(t_{a})=y_{S}$
$Y^{W}(t)$ $=m_{Y^{W}}(\tau,y_{W})+\int_{t_{a}}^{t}\sigma_{W}e^{-\lambda_{W}(t-r)}\mathrm{d}B^{W}(r),$ (114)
$Y^{S}(t)$ $=m_{Y^{S}}(\tau,y_{W},y_{S})+\int_{t_{a}}^{t}\sigma_{S}e^{-\lambda_{S}(t-r)}\mathrm{d}B^{S}(r)-\tfrac{\lambda_{S}c_{W}}{\lambda_{S}-\lambda_{W}}\sigma_{W}\int_{t_{a}}^{t}(e^{-\lambda_{W}(t-r)}-e^{-\lambda_{S}(t-r)})\mathrm{d}B^{W}(r).$ (115)
Building the linear combination leads to
$aY^{W}(t)+bY^{S}(t)=am_{Y^{W}}(\tau,y_{W})+bm_{Y^{S}}(\tau,y_{W},y_{S})+M^{W}(t)+M^{S}(t)$ (116)
@@
$M^{W}(t)$ $=\sigma_{W}a\int_{t_{a}}^{t}e^{-\lambda_{W}(t-r)}-b\tfrac{\lambda_{S}c_{W}}{\lambda_{S}-\lambda_{W}}\int_{t_{a}}^{t}(e^{-\lambda_{W}(t-r)}-e^{-\lambda_{S}(t-r)})\mathrm{d}B^{W}(r),$ (117)
$M^{S}(t)$ $=b\sigma_{S}\int_{t_{a}}^{t}e^{-\lambda_{S}(t-r)}\mathrm{d}B^{S}(r).$ (118)
Note that $M^{W}(t)$ and $M^{S}(t)$ are independent normal distributed random variables, and so is their sum. Thus, $aY^{W}(t)+bY^{S}(t)$ is a normal distributed random variable, implying that ($Y^{W}(t),Y^{S}(t))$ is bivariate Gaussian.
@@
Here $\Sigma_{WS}(\tau)$ describes the covariance between $Y^{W}$ and $Y^{S}$ defined by
$\Sigma_{WS}(\tau)$ $=\operatorname*{\textrm{Cov}}(Y^{W}(t),Y^{S}(t))=\mathbb{E}[Y^{W}(t)Y^{S}(t)]-\mathbb{E}[Y^{W}(t)]\mathbb{E}[Y^{S}(t)]$ (120)

### v3

@@ Appendix B Model Parameters
$c_{\text{p,s}}$ Storage specific heat capacity 1.025 [kJ/kgK]
$c_{\text{p,f}}$ Thermal oil specific heat capacity 2.314 [kJ/kgK]
@@
${\lambda}_{W}$ Estimate for mean reversion speed of $Y^{W}$ 0.1702 [1/$s$]
[…]
@@
${\lambda}_{S}$ Estimate for mean reversion speed of $Y^{S}$ 0.2534 [1/$s$]
@@
$n_{A}$ Number of grid-points for control AA $31$ [-]
@@
$k^{\text{Ref}}$ $\sigma$-rule parameter for $\pazocal{X}^{\text{Ref}}_{n}$ 33 [-]
$k^{\text{Ext}}$ $\sigma$-rule parameter for $\pazocal{X}_{n}$ 44 [-]
@@
$\varepsilon_{0}$ Exploration rate linearly decaying to zero 11 [-]
@@
Other Value Unit
@@
$s_{\text{Liq}}$ Liquidation price 0 [€/Mwh]
$\zeta$ Selling model parameter 0 [-]
$t_{\text{E}}$ Time horizon $120$ [h]
NN Time horizon partition $120$ [-]
$\Delta t$ Time step size 11 [h]
Appendix C Surrogate Models
As described in Walden2023 , the part-load behavior of the HTHP and the SG is represented by polynomial surrogate models, which are created using process simulation software.
The HTHP surrogate models are given by the multivariate polynomials $F_{1}$ and $F_{2}$. Here, $F_{1}$ describes the HTHX outlet temperature $T^{\text{HT,out}}$ and $F_{2}$ the consumed electrical power $P^{\text{H}}$ at a given HTHX inlet temperature $T^{\text{HT,in}}=\tau^{H}$, mass flow $\dot{m}$, LTHX air inlet temperature $T^{\text{LT,in}}=\tau^{L}$ and compressor shaft speed $d$ by
$\begin{split}T^{\text{HT,out}}&=F_{1}(\tau^{H},\dot{m},\tau^{L},d)=95.9612+0.93433\tau^{H}-0.327753\dot{m}+0.0146542\tau^{L}\\ &\hskip 8.5359pt-271.354d+0.00104853(\tau^{H})^{2}+0.0211819\tau^{H}\dot{m}-0.706122\tau^{H}d\\ &\hskip 8.5359pt-0.00388073\dot{m}\tau^{L}+0.0595068\tau^{L}d-29.4801\dot{m}d+1.04924\dot{m}^{2}+562.428d^{2}\\ &\hskip 8.5359pt-0.000716825(\tau^{H})^{2}d-2.18172\dot{m}d^{2}-151.476d^{3}+0.0229386\tau^{H}\dot{m}d\\ &\hskip 8.5359pt+0.881391\dot{m}^{2}d+0.203578\tau^{H}d^{2}-0.0405702\dot{m}^{3}-0.00148575\tau^{H}\dot{m}^{2},\\[8.61108pt] P^{\text{H}}&=F_{2}(\tau^{H},\dot{m},\tau^{L},d)=127.87+2.06342\tau^{H}+2.55723\dot{m}+0.756419\tau^{L}\\ &\hskip 8.5359pt-1164.84d-1.3829\tau^{L}d-0.0168942\tau^{H}\dot{m}-2.60579\tau^{H}d-0.540713\dot{m}^{2}\\ &\hskip 8.5359pt+13.3204\dot{m}d+1556.66d^{2},\end{split}$ (83)
The surrogate models for SG inlet and outlet temperatures as functions of mass flow read as
@@ Appendix D Details on Thermal Energy Storage Operational Modes
Due to $T^{\text{HT,out}}_{n}>T^{\text{SG,in}}$ and $l^{\text{C}}\in(0,1]$ it follows directly $T^{\text{C}}(t)\leq T^{\text{SG,in}}$, so that the charging outlet temperature never exceeds the SG inlet temperature. By charging the TES, its temperature $R$ increases linearly over time (see (3) and (5)) and based on (85), $T^{\text{C}}$ also changes with time. For this reason, the charging factor must be adjusted over time to ensure a constant SG inlet temperature. Combining of (8) with (85) and (86) leads to the following relation for the charging factor
@@
which must hold in each period $t_{n},t_{n+1})$ and describes the dependence of $l^{\text{C}}$ on the chosen control $A_{n}$ and the TES temperature $R$.
@@
Analogously to (85), the discharging outlet temperature $T^{\text{D}}$ is given by
@@
As above, the discharging factor must also be continuously adjusted in each period to ensure the constant HTHX inlet temperature $T^{\text{HT,in}}_{n}$. Substituting (88) into (89) and using (8), yields for $t\in[t_{n},t_{n+1})$
@@
Note that the time-varying charging and discharging factors (87) and (90) depending on $R$ and chosen $A_{n}$ are used to derive the control conditions in Section E.
Appendix E Details on Control Constraints
Here, we provide the derivation of the control constraints given in Subsection 3.3.
Upper Bound
A first upper bound for values of the control process $A$ is obtained from a natural upper bound of the HTHX outlet temperature $T^{\text{HT,out}}_{n}$. This follows from relation (10) and its monotonicity with respect to the compressor shaft speed. Replacing $D_{n}$ by the maximum shaft speed $d_{\text{max}}$ of the HTHP and knowing that $T^{\text{HT,in}}_{n}=T^{\text{SG,out}}$, during the charging process, it holds $T^{\text{HT,out}}_{n}\leq\tau^{\text{HT,out}}_{\text{max}}:=F_{1}(T^{\text{SG,out}},\dot{m},T^{\text{LT,in}},d_{\text{max}})$ and relation (9) implies
A second upper bound follows from the fact that the charging factor $l^{\text{C}}$ is restricted to values in $[0,1]$. During charging operation, the average TES temperature increases, i.e., $R_{n}\leq R(t)$ for $t\in[t_{n},t_{n+1}]$. Since $T^{\text{HT,out}}_{n}$ and $T^{\text{SG,in}}$ are constants, (87) implies that the charging factor $l^{\text{C}}$ also increases. Thus, to ensure that $l^{\text{C}}\in[0,1]$ on $t_{n},t_{n+1}]$, it is sufficient to require $l^{\text{C}}(t_{n+1})\leq 1$. This leads to the following upper bound for $A$ by inserting (3)-(8) into (87)
where $\zeta=n_{\text{H}}\Delta t\tfrac{\dot{m}c_{F}}{m_{S}c_{S}}$, which depends on the TES temperature $R_{n}$. In addition, the condition $R(t)\leq r_{\text{max}}=T^{\text{SG,in}}$ on $0,t_{\text{E}}]$ must also be fulfilled. During charging, $R$ increases, so it is sufficient to require $R_{n+1}=R(t_{n+1})\leq T^{\text{SG,in}}$. However, if we substitute $\overline{a}^{2}$ into (3), then this condition is satisfied and no additional constraint is applied. In summary, we obtain the following state-dependent upper bound of the control
$\overline{a}(R_{n})=\min\{\overline{a}^{1},\overline{a}^{2}(R_{n})\}.$ (93)
Lower Bound
A first lower bound results from technical reasons, namely that the HTF temperature entering the heat pump must be bounded from above, by $\tau^{\text{HT,in}}_{\text{max}}$. During discharging it holds $T^{\text{HT,out}}_{n}=T^{\text{SG,in}}$ so that we obtain from relation (9) the following:
A second lower bound follows, similar to the upper bound $\overline{a}^{2}$, from the fact that the discharging factor $l^{\text{D}}$ only takes values in $[0,1]$. During discharging, the TES temperature decreases with time, i.e., $R(t_{n+1})\leq R(t)$ for $t\in[t_{n},t_{n+1}]$. Since $T^{\text{HT,in}}_{n}$ and $T^{\text{SG,out}}$ are constants, (90) implies that the discharging factor $l^{\text{D}}$ increases. To ensure that $l^{\text{D}}\in[0,1]$ on $t_{n},t_{n+1}]$, it is therefore sufficient to require $l^{\text{D}}(t_{n+1})\leq 1$. Inserting (3)-(8) into (90), the lower bound depending on $R_{n}$ is obtained as follows
Finally, the condition $R(t)\geq r_{\text{min}}=T^{\text{SG,out}}$ must be fulfilled on $t\in[0,t_{\text{E}}]$. In discharging mode, $R$ is decreasing, so it is sufficient to require $R_{n+1}=R(t_{n+1})\geq T^{\text{SG,out}}$. Inserting $\underline{a}^{2}$ into (3) implies this condition. To summarize, the following state-dependent lower bound of the control can be specified by
$\underline{a}(R_{n})=\max\{\underline{a}^{1},\underline{a}^{2}(R_{n})\}.$ (96)
Appendix F Proofs
F.1 Proof of Lemma 3.1
@@ Appendix F Proofs
First we consider the Ornstein-Uhlenbeck process $Y^{W}(t)$ and find a solution on $t_{a},t_{b}]$ with initial condition $Y^{W}(t_{a})=y_{W}\in\mathbb{R}$. Applying Itô’s formula with $F(t,Y^{W}(t))=e^{\lambda_{W}t}Y^{W}(t)$ yields
@@
with $I_{W}(t)=\int_{t_{a}}^{t}\sigma_{W}e^{-\lambda_{W}(t-r)}\mathrm{d}B^{W}(r)$. The stochastic integral appearing in (98) is a mean-zero martingale. Using this fact, the conditional expected value of $Y^{W}(t)$ given $Y^{W}(t_{a})=y_{W}$ reads
@@
with $\tau=t-t_{a}$. The variance follows from the Itô isometry, leading to
@@
To obtain a solution for $Y^{S}(t)$, Itô’s formula is applied with $F(t,Y^{S}(t))=e^{\lambda_{S}t}Y^{S}(t)$ that gives
@@
Analogously, using the integral form with initial condition $Y^{S}(t_{a})=y_{S}$ on $t_{a},t_{b}]$ and multiplying by $e^{-\lambda_{S}t}$ yields
@@
with $I_{S,1}(t)=\int_{t_{a}}^{t}\sigma_{S}e^{-\lambda_{S}(t-r)}\mathrm{d}B^{S}(r)$ and $I_{S,2}(t)=-\lambda_{S}c_{W}\int_{t_{a}}^{t}e^{-\lambda_{S}(t-r)}Y^{W}(r)\mathrm{d}r$. The integral expression $I_{S,2}(t)$ can further be simplified by using the solution for $Y^{W}(t)$, see (98),
$I_{S,2}(t)$ $=-\lambda_{S}c_{W}\int_{t_{a}}^{t}e^{-\lambda_{S}(t-r)}\bigg(e^{-\lambda_{W}(r-t_{a})}y_{W}+\int_{t_{a}}^{r}\sigma_{W}e^{-\lambda_{W}(r-u)}\mathrm{d}B^{W}(u)\bigg)\mathrm{d}r$
@@
with $I_{S,2.1}(t)=\int_{t_{a}}^{t}e^{-\lambda_{S}(t-r)}e^{-\lambda_{W}(r-t_{a})}y_{W}\mathrm{d}r$ and $I_{S,2.2}(t)=\int_{t_{n}}^{t}\int_{t_{a}}^{r}\sigma_{W}e^{-\lambda_{S}(t-r)}e^{-\lambda_{W}(r-u)}\mathrm{d}B^{W}(u)\,\mathrm{d}r$. Evaluating each integral individually gives
$I_{S,2.1}(t)$ $=e^{-\lambda_{S}t}e^{-\lambda_{W}t_{a}}y_{W}\int_{t_{a}}^{t}e^{(\lambda_{S}-\lambda_{W})r}\mathrm{d}r=\frac{e^{-\lambda_{S}t}e^{-\lambda_{W}t_{a}}y_{W}}{\lambda_{S}-\lambda_{W}}(e^{(\lambda_{S}-\lambda_{W})t}-e^{(\lambda_{S}-\lambda_{W})t_{a}})$
$=\frac{y_{W}}{\lambda_{S}-\lambda_{W}}(e^{-\lambda_{W}(t-t_{a})}-e^{-\lambda_{S}(t-t_{a})}),$ (104)
and by interchanging the order of integration
$I_{S,2.2}(t)$ $=\int_{t_{a}}^{t}\int_{u}^{t}\sigma_{W}e^{-\lambda_{S}(t-r)}e^{-\lambda_{W}(r-u)}\mathrm{d}r\,\mathrm{d}B^{W}(u)=\sigma_{W}e^{-\lambda_{S}t}\int_{t_{a}}^{t}e^{\lambda_{W}u}\int_{u}^{t}e^{(\lambda_{S}-\lambda_{W})r}\mathrm{d}r\,\mathrm{d}B^{W}(u)$
$=\frac{\sigma_{W}e^{-\lambda_{S}t}}{\lambda_{S}-\lambda_{W}}\int_{t_{a}}^{t}e^{\lambda_{W}u}(e^{(\lambda_{S}-\lambda_{W})t}-e^{(\lambda_{S}-\lambda_{W})u})\mathrm{d}B^{W}(u)$
@@
$\begin{split}Y^{S}(t)&=y_{S}e^{-\lambda_{S}(t-t_{a})}+\int_{t_{a}}^{t}\sigma_{S}e^{-\lambda_{S}(t-r)}\mathrm{d}B^{S}(r)\\ &\hskip 8.5359pt-\frac{\lambda_{S}c_{W}}{\lambda_{S}-\lambda_{W}}\bigg(y_{W}(e^{-\lambda_{W}(t-t_{a})}-e^{-\lambda_{S}(t-t_{a})})\\ &\hskip 8.5359pt+\sigma_{W}\int_{t_{a}}^{t}(e^{-\lambda_{W}(t-r)}-e^{-\lambda_{S}(t-r)})\mathrm{d}B^{W}(r)\bigg).\end{split}$ (106)
Analogous to the calculation for $Y^{W}(t)$ it is pointed out that both stochastic integrals are martingales with mean zero Gaussian processes. Furthermore, for the expected value of $Y^{S}(t)$, we get
@@
with $\tau=t-t_{a}$. For the variance, only the variances of the stochastic integrals need to consider. Due to the independence of $B^{W}$ and $B^{S}$, this results in
$\textrm{Var}(Y^{S}(t))$ $=\mathbb{E}\bigg[\bigg(\int_{t_{a}}^{t}\sigma_{S}e^{-\lambda_{S}(t-r)}\mathrm{d}B^{S}(r)\bigg)^{2}\bigg]+\mathbb{E}\bigg[\bigg(\int_{t_{a}}^{t}\tfrac{\lambda_{S}c_{W}\sigma_{W}}{\lambda_{S}-\lambda_{W}}(e^{-\lambda_{W}(t-r)}-e^{-\lambda_{S}(t-r)})\mathrm{d}B^{W}(r)\bigg)^{2}\bigg]$
@@
The latter follows from the Itô isometry. Note that the first expected value is derived the same way as for $Y^{W}(t)$, i.e.,
@@
The second expected value is calculated by multiplying with $(e^{-\lambda_{W}(t-r)}-e^{-\lambda_{S}(t-r)})^{2}$ such that
$\mathbb{E}[I_{S,2}(t)^{2}]$ $=\bigg(\tfrac{\lambda_{S}c_{W}\sigma_{W}}{\lambda_{S}-\lambda_{W}}\bigg)^{2}\mathbb{E}\bigg[\int_{t_{a}}^{t}\big(e^{-2\lambda_{W}(t-r)}-2e^{-\lambda_{W}(t-r)}e^{-\lambda_{S}(t-r)}+e^{-2\lambda_{S}(t-r)}\big)\mathrm{d}r\bigg]$
$=\bigg(\tfrac{\lambda_{S}c_{W}\sigma_{W}}{\lambda_{S}-\lambda_{W}}\bigg)^{2}\bigg(\int_{t_{a}}^{t}e^{-2\lambda_{W}(t-r)}\mathrm{d}r-2\int_{t_{a}}^{t}e^{-(\lambda_{S}+\lambda_{W})(t-r)}\mathrm{d}r+\int_{t_{a}}^{t}e^{-2\lambda_{S}(t-r)}\mathrm{d}r\bigg)$
$=\bigg(\tfrac{\lambda_{S}c_{W}\sigma_{W}}{\lambda_{S}-\lambda_{W}}\bigg)^{2}\bigg(\frac{\Sigma_{Y^{W}}^{2}(\tau)}{\sigma_{W}^{2}}+\frac{\Sigma_{Y^{S}}^{2}(\tau)}{\sigma_{S}^{2}}-\frac{2}{\lambda_{S}+\lambda_{W}}\big(1-e^{-(\lambda_{S}+\lambda_{W})\tau}\big)\bigg)$
@@
To summarize, the variance of $Y^{S}(t)$ is given by
$\Sigma_{S}^{2}(\tau)$ $=\textrm{Var}(Y^{S}(t))$
$=\Sigma_{Y^{S}}^{2}(\tau)+\bigg(\tfrac{\lambda_{S}c_{W}\sigma_{W}}{\lambda_{S}-\lambda_{W}}\bigg)^{2}\bigg(\Sigma_{Y^{W}}^{2}(\tau)+\frac{\sigma_{W}^{2}}{\sigma_{S}^{2}}\Sigma_{Y^{S}}^{2}(\tau)-\frac{2\sigma_{W}^{2}}{\lambda_{S}+\lambda_{W}}(1-e^{-(\lambda_{S}+\lambda_{W})\tau})\bigg).$ (111)
Because $Y^{S}(t)$ is a sum of independent normal distributed random variables, it holds that
@@
In order to show that $Y^{W}(t)$ and $Y^{S}(t)$ define a bivariate Gaussian random variable, one has to shown that $aY^{W}(t)+bY^{S}(t)$ is normally distributed for all $a,b\in\mathbb{R}$. Let $a,b\in\mathbb{R}$. Let $t\in[t_{a},t_{b}]$, then by the closed-form solutions given $Y^{W}(t_{a})=y_{W}$ and $Y^{S}(t_{a})=y_{S}$, we get the linear combination
$aY^{W}(t)+bY^{S}(t)=am_{Y^{W}}(\tau,y_{W})+bm_{Y^{S}}(\tau,y_{W},y_{S})+M^{W}(t)+M^{S}(t),$ (113)
@@
$\begin{split}M^{W}(t)&=\sigma_{W}a\int_{t_{a}}^{t}e^{-\lambda_{W}(t-r)}-b\tfrac{\lambda_{S}c_{W}}{\lambda_{S}-\lambda_{W}}\int_{t_{a}}^{t}(e^{-\lambda_{W}(t-r)}-e^{-\lambda_{S}(t-r)})\mathrm{d}B^{W}(r),\\ M^{S}(t)&=b\sigma_{S}\int_{t_{a}}^{t}e^{-\lambda_{S}(t-r)}\mathrm{d}B^{S}(r).\end{split}$ (114)
Note that $M^{W}(t)$ and $M^{S}(t)$ are independent normal distributed random variables. Thus, $aY^{W}(t)+bY^{S}(t)$ is a normal distributed random variable, implying that ($Y^{W}(t),Y^{S}(t))$ is bivariate Gaussian.
@@
Here, $\Sigma_{WS}(\tau)$ describes the covariance between $Y^{W}$ and $Y^{S}$ and is defined by
$\Sigma_{WS}(\tau)$ $=\operatorname*{\textrm{Cov}}(Y^{W}(t),Y^{S}(t))=\mathbb{E}[Y^{W}(t)Y^{S}(t)]-\mathbb{E}[Y^{W}(t)]\mathbb{E}[Y^{S}(t)]$